### v2

# Digital Linearizer Based on 1-Bit Quantizations

Abstract
This paper introduces a novel low-complexity memoryless linearizer for suppression of distortion in analog frontends. It is based on our recently introduced linearizer which is inspired by neural networks, but with orders-of-magnitude lower complexity than conventional neural-networks considered in this context, and it can also outperform the conventional parallel memoryless Hammerstein linearizer. Further, it can be designed through matrix inversion and thereby the costly and time consuming numerical optimization traditionally used when training neural networks is avoided. The linearizer proposed in this paper is different in that it uses 1-bit quantizations as nonlinear activation functions and different bias values. These features enable a look-up table implementation which eliminates all but one of the multiplications and additions required for the linearization. Extensive simulations and comparisons are included in the paper, for distorted multi-tone signals and bandpass filtered white noise, which demonstrate the efficacy of the proposed linearizer.
I Introduction
Digital linearization (post-correction and pre-distortion) are used for suppressing nonlinearities (distortion) emanating from imperfections in analog circuits and systems. In this paper, the focus is on digital linearization needed in e.g. analog front-ends on the receiver side in communication systems. The nonlinearities emanate from the analog-to-digital converters (ADCs) and its preceding components including filters, amplifiers, and mixers. In addition to suppressing the nonlinearities, it is important to develop low-complexity digital linearizers to enable energy-efficient hardware implementations, especially in systems with very high data rates. Efficient digital linearizers also enable the use of ADCs with relaxed linearity requirements and resolutions (fewer bits) which can reduce their energy consumption substantially [1].
This paper concerns frequency-independent distortion for which one can use memoryless polynomial modeling and linearization [^1] Memoryless nonlinearity modeling and linearization is typically sufficient for narrow to medium analog bandwidths and resolutions. To reach higher resolutions over wider frequency bands, one may need to incorporate memory (subfilters) in the modeling and linearization.. Recently, [2] introduced a low-complexity memoryless linearizer (Fig. 2 in Section III) that is inspired by neural networks, but has orders-of-magnitude lower complexity than conventional neural-network schemes considered in this context [3, 4], and it can be designed through matrix inversion, thereby avoiding the costly and time-consuming numerical optimization that is traditionally used when training neural networks. It was also demonstrated in [2] that it can outperform the conventional parallel memoryless Hammerstein linearizer [5] (Fig. 1 in Section II). Nevertheless, the linearizer introduced in [2] still requires several multiplications and additions per linearized sample.
In this paper, a new linearizer is proposed which enables a look-up table implementation, which eliminates all but one of the multiplications and additions required for the linearization. The proposed linearizer is based on the one in [2] but differs in two ways. Firstly, the nonlinear activation functions are here 1-bit quantizers instead of the rectified linear unit (ReLU) or modulus operations adopted in [2]. Secondly, instead of determining the bias values in the design as in [2], they are here a-priori selected carefully. An additional advantage of the proposed linearizer, over that in [2], is that it requires only one design (one matrix inversion). In [2], several designs are used to find the optimal bias values.
Following this introduction, Section II briefly recapitulates the linearization problem and reviews the Hammerstein linearizer and the one in [2]. Section III considers the proposed linearizer and its design. Section IV provides evaluations and comparisons, demonstrating the efficacy of the proposed linearizer. Finally, Section V concludes the paper.
II Distortion and Linearization
Consider a desired discrete-time signal $x(n)=x_{a}(nT)$, representing a sampled version of an analog signal $x_{a}(t)$ with a uniform sampling interval $T$. In practice, the output of an ADC will not be $x(n)$ but a distorted version of it, say $v(n)$. In this paper, it is assumed that the distortion is memoryless, in which case $v(n)$ is commonly modeled as a memoryless polynomial according to [5]
$v(n)=a_{0}+a_{1}x(n)+\sum_{p=2}^{P}a_{p}x^{p}(n),$ (1)
where $a_{0}$ is a constant (offset), $a_{1}$ is a linear-distortion coefficient and $a_{p}$, $p=2,3,\ldots,P$, are nonlinear-distortion coefficients. Additionally, the signal contains quantization noise, but it is here excluded from the mathematical expressions for the sake of simplicity. Before proceeding, it is also stressed that the proposed linearizer (Section III) as well as the existing ones to be reviewed below do not require that the distorted signal is in the form of (1). In the simulations in this paper, the model in (1) is used for generating sets of test and evaluation signals in order to assess the performance of the different linearizers.
Figure 1: Hammerstein linearizer with the upper (lower) dashed box indicating the linear branch (nonlinear branches).
Given the distorted signal $v(n)$, the linearization amounts to generating a compensated signal, say $y(n)$, in which the distortion has been suppressed (ideally removed). In the traditional Hammerstein linearizer, illustrated in Fig. 1, $y(n)$ is generated as
$y(n)=d_{0}+d_{1}v(n)+\sum_{k=2}^{K}d_{k}v^{k}(n).$ (2)
In total, this linearizer requires 2K12K-1 multiplications and $K$ additions per corrected output sample, with $K$ multiplications to produce $d_{k}v^{k}(n)$ and K1K-1 multiplications for all the $v^{k}(n)$ computations. Alternatively, one may use the so-called Wiener linearizer [5] where the multiplications by $d_{k}$, $k=2,3,\ldots,K$, seen in Fig. 1, are carried out before the operations $v^{k}(n)$, but it does not change the implementation complexity and performance.
Recently, the linearizer seen in Fig. 2 was introduced [2]. The compensated signal $y(n)$ is then generated as
$y(n)=c_{0}+c_{1}v(n)+\sum_{m=1}^{N}w_{m}f_{m}(v(n)+b_{m})$ (3)
with $f_{m}$, $m=1,2,\ldots,N$, representing a set of $N$ nonlinear functions. In [2], the ReLU operations $f_{m}(v(n))=\max\{0,v(n)+b_{m}\}$ or modulus operations $f_{m}(v(n)+b_{m})=|v(n)+b_{m}|$ were used due to their simplicity in hardware implementations [6]. Further, the bias values $b_{m}$ were uniformly distributed between $b_{\max}$ and $b_{\max}$, i.e., $b_{m}=-b_{\max}+2(m-1)b_{\max}/(N-1)$, where the value of $b_{\max}$ was determined in the design. It was demonstrated in [2], that the linearizer in Fig. 2 can outperform the Hammerstein linearizer even when the nonlinearities have been generated through the memoryless polynomial model in (1). However, the linearizer in Fig. 2 still requires $N+1$ multiplications and $2N+1$ additions per corrected output sample.
Figure 2: Linearizer in [2] with the upper (lower) dashed box indicating the linear branch (nonlinear branches).
III Proposed Linearizer
In this paper, a new linearizer is proposed which enables a look-up table implementation, thereby eliminating the multiplications and additions in the $N$-branch nonlinearity part of the linearizer in Fig. 2. It is based on the one in Fig. 2 but it has two distinct differences. The first one is that the nonlinear functions are here 1-bit quantizations according to
$f_{m}\big(v(n)+b_{m}\big)=\begin{cases}1,&v(n)+b_{m}\geq 0,\\ 0,&v(n)+b_{m}<0.\end{cases}$ (4)
The second difference is that, instead of determining $b_{\max}$ in the design [2], it is here selected as
$b_{\max}=\frac{N-1}{N+1}.$ (5)
The bias values $b_{m}$ are then again uniformly distributed between $b_{\max}$ and $b_{\max}$. Utilizing (5), $b_{m}$ are thus chosen as
$b_{m}=-1+\frac{2m}{N+1},\,m=1,2,\ldots,N.$ (6)
The implication of the choices in (4)–(6) is two-fold. Firstly, for each sample index $n$, since the output of each nonlinear operation $f_{m}\big(v(n)+b_{m}\big)$ is either one or zero, and $b_{m}$ are fixed and uniformly distributed between $b_{\max}$ and $b_{\max}$, there are $N+1$ possible combinations of zeros and ones in the $N$ inputs to $w_{m}$, and these combinations contain consecutive zeros followed by consecutive ones (except for two combinations with only zeros or only ones). For example, with $N=4$, the only possible combinations are $0000$, $0001$, $0011$, $0111$, and $1111$. This means that the $N$ multiplications by $w_{m}$ followed by the corresponding N1N-1 additions, and the addition of $c_{0}$, can be implemented with a look-up table of size $N+1$ and with entries $u_{q}$, $q=0,1,\ldots,N$, where
$u_{q}=\begin{cases}c_{0},&q=0,\\ c_{0}+\sum_{i=N-q+1}^{N}w_{i},&1\leq q\leq N.\end{cases}$ (7)
Secondly, with $b_{\max}$ as in (5), and with $b_{m}$ uniformly distributed between $b_{\max}$ and $b_{\max}$, it is readily shown that the whole signal-value region from 1-1 to $1$, of the input signal $v(n)$, can be divided into $N+1$ consecutive sub-regions of the same size $2/(N+1)$, with each sub-region corresponding to a unique combination of the $N+1$ possible zero/one combinations at the quantizers’ outputs. This is exemplified in Fig. 3 for $N=7$, and it means in general that neither the $N$ additions by the bias values $b_{m}$ nor the $1$-bit quantizations and subsequent multiplications have to be carried out. Instead, the look-up table can be directly addressed by identifying and mapping the input signal level to a corresponding memory address. For this, it suffices to use $\lceil\log_{2}(N+1)\rceil$ memory-address bits as exemplified in Fig. 3. The proposed linearizer can therefore be implemented as Fig. 4 depicts [^2] By scaling the linear and nonlinear branches by $1/c_{1}$, one could remove the multiplication in the linear branch. However, in a practical implementation, scaling of the signal levels is required [7], which means that one multiplication is generally needed anyhow.. It can be equivalently described by the scheme in Fig. 5, since the look-up table based implementation corresponds to adding a time-varying value $w\big(v(n)\big)$ which can take on the $N+1$ values in the table and whose value for each $n$ is determined by the value of $v(n)$. Before proceeding, it is noted that post-correction methods for ADCs utilizing look-up tables have appeared earlier, as in [8] for the memoryless case. However, those methods are different as they target small-scale nonlinearities caused by nonideal quantizers (deviations from uniform quantizations) and the correction table is then designed based on ADC output-code statistics. Further, those correction methods (with one fixed table) typically offer modest signal-to-noise-and-distortion ratio (SNDR) improvements for some frequencies but can even deteriorate the SNDR for other frequencies [8]. The linearizer in this paper targets more general nonlinearities and larger SNDR improvements, and it is designed to correct all frequencies in a prespecified frequency band. To further stress the difference, it is noted that 1) the proposed linearizer is designed via the basic structure in Fig. 2 and it can also be implemented in that way (but the look-up table implementation can be more efficient), which is not the case for the previous look-up-table-based methods, and 2) the size of the memory ($N+1$) in the proposal can be kept relatively small as it is determined by the number of nonlinear branches ($N$), and not by the number of data-bits combinations utilized when designing the tables in the previous methods.
Figure 3: Signal levels and corresponding inputs to $w_{m}$ and memory address for $N=7$.
Figure 4: Implementation of the proposed linearizer using a memory of size $N+1$ and $\lceil\log_{2}(N+1)\rceil$ memory-address bits (see Footnote 2).
Figure 5: Equivalence of the implementation in Fig. 4.
III-A Design
An advantage of the proposed linearizer, as well as that in [2], over traditional neural networks is that it can be readily designed through matrix inversion, thereby eliminating the costly and time-consuming numerical optimization that is traditionally used when training neural networks. Further, an advantage of the proposed linearizer, over that in [2], is that it requires only one design (one matrix inversion) as $b_{\text{max}}$ and thereby $b_{m}$ are predetermined according to (5) and (6). In [2], in contrast, a number of designs (matrix inversions) were carried out, one for each value of $b_{\text{max}}$ over a set of different values, and the best of the so obtained solutions was selected.
Since the bias values $b_{m}$ are predetermined in the proposed linearizer, the design amounts to determining the parameters $c_{0}$, $c_{1}$, and $w_{m}$, $m=1,2,\ldots,N$, so that the output signal $y(n)$ approximates the desired signal $x(n)$ as well as possible, here in the least-squares sense. Even if the proposed linearizer can be implemented with a look-up table, its design is most easily carried out via the basic scheme in Fig. 2 with the nonlinear functions in (4) and bias values in (5) and (6). Further, in the design, $c_{1}v(n$) in (3) is replaced with $v(n)+\Delta c_{1}v(n)$ and then $\Delta c_{1}$ is determined. Thereby, all parameters to be determined are small (zero when the distortion is zero). In addition, $L_{2}$-regularization is used to reduce the risk of ill-conditioned matrices and avoid large parameter values. Based on the above, the following design procedure is proposed.
  1. 1.
    Generate a set of $R$ reference signals $x_{r}(n)$ and their distorted signals $v_{r}(n)$, for $r=1,2,\ldots,R$, utilizing a signal model such as that in (1) or measured data.
  2. 2.
    Minimize the cost function $E$ defined by
    $E=\frac{1}{RL}\sum_{r=1}^{R}\sum_{n=1}^{L}\big(y_{r}(n)-x_{r}(n)\big)^{2},$ (8)
    where $L$ denotes the length of the signal. To solve this problem, define $\bf w$} as a $(N+2)\times 1$ column vector encompassing all coefficients $w_{m}$ for $m=1,2,\ldots,N$, $c_{1}$, and $c_{0}$, and let ${\bf{A}}_{r}$ be an $L\times(N+2)$ matrix with columns $l$ for $l=1,2,\ldots,N+2$, filled with the $L$ samples $f_{m}\big(v_{r}(n)+b_{m}\big)$ for $m=1,2,\ldots,N$, $L$ samples of $v_{r}(n)$, and $L$ ones for the constant $c_{0}$. Minimizing $E$ in (8) in the least-squares sense then gives the solution
    ${\bf w}={\bf A}^{-1}{\bf b},$ (9)
    where
    ${\bf A}=\lambda{\bf I}+\frac{1}{RL}\sum_{r=1}^{R}{{\bf{A}}_{r}^{\top}{\bf A}_{r}},\quad{\bf b}=\frac{1}{RL}\sum_{r=1}^{R}{{\bf A}_{r}^{\top}{\bf b}_{r}}.$ (10)
    Here, ${\bf{A}}_{r}^{\top}$ is the transpose of ${\bf{A}}_{r}$, ${\bf{b}}_{r}$ is an $L\times 1$ column vector with the samples $x_{r}(n)-v_{r}(n)$, and $\lambda{\bf I}$ is a diagonal matrix with small diagonal values $\lambda$ for the $L_{2}$-regularization. The linearized output (row vector) ${\bf y}_{r}=y_{r}(n)$, $n=1,2,\ldots,L$, is given by
    ${\bf{y}}_{r}={\bf{v}}_{r}+{\bf{w}}^{\top}{\bf{A}}_{r}^{\top}.$ (11)
    where ${\bf v}_{r}=v_{r}(n)$, $n=1,2,\ldots,L$ is also a row vector.
  3. 3.
    Assess the performance of the linearizer over $M$ signals, where $M\gg R$, to ensure a robust validation.
IV Evaluations and Comparisons
For the evaluations and comparisons, we assume the same distorted signal $v(n)$ as in [2] where $v(n)$ is modeled as in (1) with $a_{0}=0$, $a_{1}=1$, and $a_{p}=(-1)^{p}\times 0.15/p$, for $p=2,3,\ldots,P$ with $P=10$.
Example 1: We consider the multi-tone signal
$x(n)=G\times\sum_{k=1}^{31}A_{k}\sin(\omega_{k}n+\alpha_{k}),$ (12)
where $A_{k}=1$ for all $k$, and $\alpha_{k}$ are randomly chosen from $\{\pi/4,-\pi/4,3\pi/4,-3\pi/4\}$, which corresponds to QPSK modulation. The frequencies $\omega_{k}$ are given by
$\omega_{k}=\frac{2\pi k}{64}+\Delta\omega,$ (13)
in which case the signal corresponds to the quadrature (imaginary) part of $31$ active subcarriers in a $64$-subcarrier OFDM signal with a random frequency offset $\Delta\omega$. In the design and evaluation we use, respectively, $R=1$ and $M=2500$ signals with randomly generated frequency offsets assuming uniform distribution between $\pi/64$ and $\pi/64$, quantized to $8$ bits, and of length $L=8192$. The gain $G$ is selected so that the distorted signal is below one in magnitude. For the $L_{2}$-regularization, we use $\lambda=0.0002$. Figure 6 plots the spectrum before and after linearization for one of the signals. Figure 7 plots the mean SNDR over 2500 signals for each linearizer instance (with an SNDR variance of some 0.5 dB for all instances) versus the number of branches for the proposed linearizer, the one in [2], and the Hammerstein linearizer, designed in the same least-squares sense (for all designs, the optimized parameter values were quantized to $12$ bits). For the signals considered in this example, quantized to $8$ bits, the signal-to-noise ratio (SNR) is some $42$ dB without distortion, and the SNDR is some $25$ dB for the distorted signal before the linearization. Hence, the linearizer can at most improve the SNDR by $17$ dB in this example, which corresponds to almost $3$ bits improvement.
As seen in Fig. 7, for the proposed linearizer, the SNDR does not increase monotonically. This is because its bias values are not optimized but instead selected according to (5) and (6) to enable the look-up table implementation. Further, the proposed linearizer requires more nonlinear branches than the other two methods to reach the same SNDR. However, to assess the implementation complexity, one should not consider the number of branches but instead the complexity required to linearize each output sample. With $N$ nonlinear branches, the Hammerstein linearizer (where N=K1)N=K-1) requires $2N+1$ multiplications and $N+1$ additions whereas the linearizer in [2] requires $N+1$ multiplications and $2N+1$ additions. In general, multiplications are considerably more complex to implement than additions and the linearizer in [2] is thus more efficient than the Hammerstein linearizer for the same number of nonlinear branches. The proposed linearizer, implemented with a look-up table, requires only one multiplication and one addition per output sample. Hence, in terms of computational complexity, this linearizer clearly outperforms the other two, as illustrated in Fig. 8 which plots the number of multiplications versus the SNDR for the three methods. However, the proposed linearizer also have the additional cost of the memory look-up implementation, which is dependent upon the hardware platform.
Example 2: To further illustrate the robustness of the proposed linearizer designed in Example 1, we have also evaluated it for the same type of multi-sine signal as in Example 1 but with some of the subcarriers set to zero, and a bandpass filtered white-noise signal covering 50% of the Nyquist band. As illustrated in Figs. 9 and 10 for one of each of these signals, essentially the same result is obtained. Less than 1 dB SNDR degradation compared to the linearized signals considered in Example 1 was observed.
Figure 6: Spectrum before and after linearization for a multi-sine signal using the proposed linearizer with $N=32$ nonlinear branches.
Figure 7: SNDR versus the number of nonlinear branches.
Figure 8: SNDR versus the number of multiplications required per corrected output sample.
Figure 9: Spectrum before and after linearization for a multi-sine signal with null subcarriers using the proposed linearizer with $N=32$ nonlinear branches.
Figure 10: Spectrum before and after linearization for a bandpass filtered white-noise signal using the proposed linearizer with $N=32$ nonlinear branches.
V Conclusions
This paper introduced a low-complexity memoryless linearizer for suppression of distortion in analog frontends. It is based on our recently introduced linearizer in [2], which is inspired by neural networks but offers orders-of-magnitude lower complexity, and can outperform the conventional parallel memoryless Hammerstein linearizer. Further, it can be designed through matrix inversion and thereby the costly and time consuming numerical optimization traditionally used when training neural networks is avoided. The proposed linearizer is different from [2] in that it uses 1-bit quantizations as nonlinear activation functions (instead of ReLU and modulus operations) and different and carefully selected bias values. These features enable a look-up table implementation (Fig. 4) which means that the linearization only requires one multiplication and one addition to correct each output sample, regardless of the number of nonlinear branches ($N$).
Examples included also demonstrated that the same SNDR can be reached with the proposed linearizer, as with the previous linearizers, by increasing the number of nonlinear branches. For the same SNDR improvement, the proposed linearizer is thus superior in terms of computational complexity, but it also requires the additional memory. Hence, there is a trade-off between computational complexity and the additional cost of the memory look-up implementation. Again, it is stressed though that the size of the memory ($N+1$) in the proposal is relatively small since it is determined by the number of nonlinear branches ($N$), not by the number of data-bits combinations utilized when designing the tables in the previous look-up-table-based methods. Further studies thus include the utilization of both $1$-bit quantization and ReLU/modulus operations in the linearizer in order to find the lowest implementation complexity, given that the cost of different operations is available for a specific hardware platform. The extension to memory linearizers addressing frequency dependent nonlinearities will also be studied.
Acknowledgment
This work belongs to the project Baseband Processing for Beyond 5G Wireless (B02), financially supported by ELLIIT.